OpenStudy (rissaunnie):

Simplify this expression. 3x^2(7x^2 - 9)

OpenStudy (welshfella):

3x^2 * 7x^2 + 3x^2*-9

OpenStudy (anonymous):

3x2(7x2-9) Final result : 3x2 • (7x2 - 9) Reformatting the input : Changes made to your input should not affect the solution: (1): "x2" was replaced by "x^2". 1 more similar replacement(s). Step by step solution : Step 1 : Raise x to the 2nd power Exponentiation : Equation at the end of step 1 : (3 • (x2)) • ((7 • x2) - 9) Step 2 : Raise x to the 2nd power Exponentiation : Equation at the end of step 2 : (3 • x2) • (7x2 - 9) Step 3 : Multiply 3x2 by 7x2-9 Trying to factor as a Difference of Squares : 3.1 Factoring: 7x2-9 Theory : A difference of two perfect squares, A2 - B2 can be factored into (A+B) • (A-B) Proof : (A+B) • (A-B) = A2 - AB + BA - B2 = A2 - AB + AB - B2 = A2 - B2 Note : AB = BA is the commutative property of multiplication. Note : - AB + AB equals zero and is therefore eliminated from the expression. Check : 7 is not a square !! Ruling : Binomial can not be factored as the difference of two perfect squares Final result : 3x2 • (7x2 - 9)

OpenStudy (rissaunnie):

21x2 - 9 21x2 - 27x2 21x4 - 9 21x4 - 27

OpenStudy (anonymous):

welsha will you go to my question i taged you in

OpenStudy (rissaunnie):

that is the answer options